A stone of mass 1kg is thrown at 10m/s upward Making an angle 37° with the horizontal from a building thta is 20m high.using law of conservation of energy . calculate the speed when the stone hits the ground.


1
Expert's answer
2021-02-15T17:40:53-0500

By the law of energy conservation :

"\frac{mv^2}{2}+mgh=\frac{mu^2}{2}"

where, "v" is inutial speed, "u" is final speed, "h" is building height

When the stone falls back down and reaches the horizontal position again, it will have the same angle as at launch: "37\degree" .

Horizontal component of velocity:

"v_x=u_x=10cos37\degree=8" m/s; this does not change

Vertical component of initial velocity:

"v_y=10sin37\degree=6" m/s

So:

"u_y=\sqrt{v^2_y+2gh}=\sqrt{6^2+2\cdot9.8\cdot20}=20.7" m/s

The final speed:

"u=\sqrt{8^2+20.7^2}=22.2" m/s